OpenStudy (anonymous):

I will give medal and be a fan please I really need help I have a test tomorrow and I need to learn it so please explain this to be thank you. I need this in matrix way till the end. Solve the system of equations using Gaussian elimination or gauss-Jordan elimination. 5x+4y+z=3 4x-3y-z=37 5x+y+2z=16

OpenStudy (anonymous):

Can you please help me @twvogels

OpenStudy (amistre64):

do you know how to do elimination?

OpenStudy (anonymous):

I do I know the answer but I need to learn it in matrix

OpenStudy (amistre64):

if you can do elimination, then you are doing the matrix

OpenStudy (anonymous):

Ok this the example my teacher give and I don't know this at all

OpenStudy (amistre64):

5x+4y+z=3 4x-3y-z=37 5x+y+2z=16 use the first equation to eliminate the x parts of the next 2 equations

OpenStudy (amistre64):

"using a matrix" does not change the math that goes on, nor does it make it any simpler to do by hand. so its best to see it done in its equation form to understand that its not anything magical happening

OpenStudy (anonymous):

I know but that's the way I need to do it on my test

OpenStudy (amistre64):

its the same process, all that changes is how you format the information ...

OpenStudy (amistre64):

a matrix just 'removes' the variables so all that you work on is the numbers attached to them ... which is all you work on in any case.

OpenStudy (amistre64):

equation format: 5x+4y+z=3 4x-3y-z=37 5x+y+2z=16 ------------------ matrix format: 5 4 1 3 4 -3 -1 37 5 1 2 16 ---------------- use the first equation/row to eliminate the x part of the other 2

OpenStudy (amistre64):

5x +16/4y+ 4/4z = 12/4 -5x+15/4 y +5/4z = -5(37)/4 ------------------------- 0x +31/4y +9/4z = -173/4 5x +4y+ z = 3 -5x -y -2z = -16 ------------------ 0x +3y -z = -13 ------------------------- 5x +4y + z = 3 0x +31y +9z = -173 0x +3y -z = -13 now use the second equation to eliminate the y parts of the other 2 ....

OpenStudy (amistre64):

its the same process in matrix form ... there is no new mathing that is happening.